a 6000 seat theater has tickest for sale at $27 and$40. how many tickets should be sold at each price for a sellout preformance to genorate a total revenue of $181,500?
Answer by Maths68(1474)

a 6000 seat theater has tickest for sale at $27 and$40. how many tickets should be sold at each price for a sellout preformance to genorate a total revenue of $181,500?
Let
$27 tickets sold = x
$40 tickets sold = 6000-x
Total Amount = 181500

$27 tickets sold+$40 tickets sold = Total Amount
27*x+40(6000-x)=181500
27x+240000-40x=181500
27x-40x=181500-240000
-13x=-58500
-13x/-13=-58500/-13
x=4500

$27 tickets sold = x = 4500
$40 tickets sold = 6000-x = 6000-4500=1500